If your mean is 33.7 and your standard deviation is 33.7 what is your z score?
Z-score is the x value minus the mean, all divided by the standard deviation; or z=(x-mu)/sigma. The "x" value needs to be given to answer the question.
How many possible hands are there for only getting 3 cards that make a straight flush?
48. Don't know if K A 2 x 4 counts but if it does then it is 52.
A 2 3 x 4
2 3 4 x 4
3 4 5 x 4
4 5 6 x 4
5 6 7 x 4
6 7 8 x 4
7 8 9 x 4
8 9 10 x 4
9 10 J x 4
10 J Q x 4
J Q K x 4
Q K A x 4
What is the probability that all 4 cards will be the same suit?
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
A claim's initial plausibility is assessed by what?
It is generally reasonable to accept the claim if it does not conflict with our store of background information and if it comes from a reliable source.
A claim's initial plausibility is a measure of how well it "squares" with our background information and our own observations.
The chance of rolling a specific number on a six-sided die is 1/6. The same 1/6 chance applies to the second die.
Multiply them together to find the chance of both events occurring: 1/6 x 1/6 = 1/36, or about 2.78%
What is the probability of getting all heads on a toss of three fair coins?
1/2 * 1/2 * 1/2 = 1/8 = 12.5%
What are the formulas in probability?
There are many, many formulae:
for different probability distribution functions,
for cumulative distribution functions,
for moment generating functions,
for means, variances, skewness, kurtosis and higher moments.
There are many, many formulae:
for different probability distribution functions,
for cumulative distribution functions,
for moment generating functions,
for means, variances, skewness, kurtosis and higher moments.
There are many, many formulae:
for different probability distribution functions,
for cumulative distribution functions,
for moment generating functions,
for means, variances, skewness, kurtosis and higher moments.
There are many, many formulae:
for different probability distribution functions,
for cumulative distribution functions,
for moment generating functions,
for means, variances, skewness, kurtosis and higher moments.
The factors of 20 are 1, 2, 4, 5, 10, and 20. We can take 10 and 20 off the list, because they're not on the six-sided dice.
So, 4 out of 6 sides of the dice will give us a factor of 20. That means we can expect (if the dice is fair) to get a factor of 20 about 4/6 of the time. 4/6 simplified is 2/3.
What is a systematic way to find all possible outcomes?
The answer is not easy: it depends on the situation that you are studying. It is simple to identify all the possible outcomes of rolling a die and tossing a coin: it is the set of ordered pairs defined as the set
{(n, a), where n is an integer from 1 to 6 and a is H or T}.
In general, it is not at all simple to determine all possible outcomes. There may be outcomes that you did not expect or and make allowance for. In practise, therefore, it is sometimes best to have a category for "other outcomes".
In principal, a tossed coin could end up on its edge, or at least, end up leaning against some obstruction so that it is not flat on H or T.
How long can you be trainee for?
Technically indefinitely, only once your superior is satisfied with your progress would no longer be considered a trainee.
Give 5 solving problems about partitive proportions?
tanong mo nalang sa guro mong tanga! bakit hindi mo alam? hindi ka ba tinuturuan ng guro mo? oh sadya lang bobo yung guro mo?
How do you solve a mean squared deviation?
You cannot "solve" a mean squared deviation". You can calculate it or use it, but there is nothing to solve!
How many numbers of ways you can arrange 32 different cards between 4 players without repetition?
I'm pretty sure the answer is 32!/(8!)4 which is 99 561 092 450 391 000.
The result of tossing the coin would not affect which number was selected. So we say that these two events are independent. We can therefore assess the probability of each of them separately and then multiply the two probabilities together for a final result.
Probability of getting tails: 1/2 (since there is one way of getting heads out of two possibilities)
Probability of getting zero: 1/10 (since there is one way of getting zero out of ten possibilities)
Overall probability: 1/2 x 1/20 = 1/20
What is the difference between a bias and a random error?
Bias is systematic error. Random error is not.
A probability based upon an event that has already occurred?
May - or may not - be a conditional probability. A conditional probability is not becessarily chronologically structured.
What is the probability of rolling more than 8 with one roll of the dice?
The probability of rolling a nine or higher with one pair of dice is 10 in 36, or 5 in 18, or about 0.2778.
Out of the 36 possible permutations, the 10 that match the criteria are (6-3) (5-4) (4-5) (3-6) (6-4) (5-5) (4-6) (6-5) (5-6) (6-6).